# SGD with Clipping is Secretly Estimating the Median Gradient

### v1

@@ 3 Warm-up: Sample Median Gradient
Figure 1: First and second moment of sample median/mean for standard ($\alpha=1.1$)-stable distribution. We approximate $\mathbb{E}$ by running $1000$ trials. Even for small sample size, the sample median has small variance (cf. (7)), in contrast to the sample mean.
4 Iterative Methods for Median Estimation
We now focus on the following gradient estimation problem: let $\mathcal{G}$ denote the (fixed) distribution over noisy gradients $\bm{g}$, and consider the problem of approximating
In the previous section, we used the sample median. It follows that solving (8) with $\mathcal{D}=\|\cdot\|_{p}$, and with $\mathcal{G}$ being the empirical distribution over a sample $(\bm{g}^{(1)},\ldots,\bm{g}^{(n)})$, the solution to (8) is the sample median (3).
For solving (8), we consider only stochastic iterative methods, that can be applied for any distribution $\mathcal{G}$. The most commonly used such method is SGD, which has been used for computing the median (Cardot et al., 2011). But here we advocate for a different method: The stochastic proximal point method (SPP(Asi & Duchi, 2019; Davis & Drusvyatskiy, 2019). In cases where SPP has a closed form update, it is often preferred to SGD since is easier to tune (Asi & Duchi, 2019; Milzarek et al., 2022). Fortunately, as we show next, the SPP method enjoys closed-form updates when $\mathcal{D}\in\{\|\cdot\|_{1},\|\cdot\|_{2},\tfrac{1}{2}\|\cdot\|^{2}\}$.
The update of SPP with stepsize $\tau>0$ applied to (8) can be written as (see details in Appendix B.3)
$\bm{m}_{t+1}\;=\;\bm{g}_{t}+\mathrm{prox}_{\tau\mathcal{D}}(\bm{m}_{t}-\bm{g}_{t}),$ (9)
$\mathrm{prox}_{\mathcal{D}}(\bm{m}):=\underset{\bm{y}\in\mathbb{R}^{d}}{\operatorname{argmin}}\;\mathcal{D}(\bm{y})+\frac{1}{2}\|\bm{y}-\bm{m}\|^{2}.$
Because we are interested in the $\ell_{p}$-median, we first characterize the SPP update (9) for when $\mathcal{D}$ is the $\ell_{p}$-norm, and show it simply require to project the sampled gradient onto a certain ball of radius $\tau$ centered at $\bm{m}_{t}$.
Proposition 4.1.
$=\mathrm{Proj}_{\mathbb{B}_{q}(\bm{m}_{t},\tau)}\big{(}\bm{g}_{t}\big{)}.$ (11)
Using Proposition 4.1 we can now develop closed form updates of the SPP method for when $\mathcal{D}$ is the $\ell_{1}$- or the $\ell_{2}$-norm, with both methods being related to clipping.
Vectorwise clipping.
If we choose $\mathcal{D}=\|\cdot\|_{2}$, SPP performs projections onto the $\ell^{2}$-ball which is also known as clipping (Khirirat et al., 2023; Koloskova et al., 2023).
Corollary 4.2.
For $\mathcal{D}=\|\cdot\|_{2}$ update (10) is given by
$\bm{m}_{t+1}$ $=\bm{m}_{t}+\mathrm{clip}_{\tau,2}(\bm{g}_{t}-\bm{m}_{t}),$ (12)
This corollary shows that the SPP method for computing the $\ell_{2}$-median can be written as $\bm{m}_{t+1}=\beta_{t}\bm{m}_{t}+(1-\beta_{t})\bm{g}_{t}$, with $\beta_{t}:=1-\tfrac{\tau}{\max\{\tau,\|\bm{g}_{t}-\bm{m}_{t}\|_{2}\}}$. In this sense, (12) is momentum where the momentum coefficient depends on how far $\bm{g}_{t}$ is from the previous estimate $\bm{m}_{t}$. This allows the iterates to be more robust to outliers, as samples which are too far from $\bm{m}_{t}$ will be down weighted (see also Fig. 2).
Componentwise clipping.
If we choose $\mathcal{D}=\|\cdot\|_{1}$, SPP performs projections onto the $\ell^{\infty}$-ball, also known as componentwise clipping. Here as well, clipping protects against large individual entries by shrinking them independently.
Corollary 4.3.
For $\mathcal{D}=\|\cdot\|_{1}$ update  (10) is given by
$\bm{m}_{t+1}=\bm{m}_{t}+\mathrm{clip}_{\tau,1}(\bm{g}_{t}-\bm{m}_{t}),$ (13)
SPP and momentum.
It is easy to see that the solution to problem (8) with $\mathcal{D}=\tfrac{1}{2}\|\cdot\|^{2}$ is the mean $\mathbb{E}\left[\bm{g}\right]$. Computing (9) for this choice of $\mathcal{D}$, we obtain
$\bm{m}_{t+1}=(1-\tfrac{\tau}{1+\tau})\bm{m}_{t}+\tfrac{\tau}{1+\tau}\bm{g}_{t}.$ (14)
Interestingly, this is the (heavy-ball) momentum gradient estimator (Polyak, 1964; Sutskever et al., 2013), where the momentum coefficient is given by $\beta=1/(1+\tau)\in[0,1).$
Rates for SPP
Now that we have presented how to compute the iterates (9) in practice, let us state its convergence properties. Since all these methods are instantiates of the SPP method, we can use standard results such as (Davis & Drusvyatskiy, 2019, Thm. 4.4).
Proposition 4.4.
Let $\mathcal{D}$ be a norm, let $\varphi(\bm{m}):=\mathbb{E}_{\bm{g}\sim\mathcal{G}}\left[\mathcal{D}(\bm{m}-\bm{g})\right]$, and let $\bm{m}_{*}\in\underset{\bm{m}\in\mathbb{R}^{d}}{\operatorname{argmin}}\;\varphi(\bm{m})$. If $\bm{m}_{t}$ is generated by (9) then for $\bar{\bm{m}}_{T}=\frac{1}{T}\sum_{t=1}^{T}\bm{m}_{t}$ we have
$\mathbb{E}\left[\varphi(\bar{\bm{m}}_{T})-\varphi(\bm{m}_{*})\right]=O\left(\frac{\|\bm{m}_{0}-\bm{m}_{*}\|^{2}}{\tau T}+\tau\right).$
In Appendix B we give further examples for $\mathcal{D}$, such as the Huber loss. We also compare SPP with SGD for solving (8), showing that sign-SGD is related to $\ell_{1}$-median estimation.
5 Gradient Estimator in the Wild
$\begin{array}[]{llr}\bm{m}_{t+1}&=\quad\bm{g}_{t}+\mathrm{prox}_{\tau\mathcal{D}}(\bm{m}_{t}-\bm{g}_{t}),\\ \bm{w}_{t+1}&=\quad\bm{w}_{t}-\eta_{t}\bm{m}_{t+1}.\\ \end{array}$ (17)
Formally, $\bm{m}_{t}$ is now an estimate of the median of the gradients across iterations, which is in contrast to (5), where $\bm{m}_{t}$ was an approximation of the sample median at fixed $\bm{w}_{t}$. With the results from Section 4, we immediately obtain the following special cases (here $p\in\{1,2\}$):
(ii)  if $\mathcal{D}=\|\cdot\|_{p}$ :  $\begin{cases}\bm{m}_{t+1}=\bm{m}_{t}+{\rm clip}_{p,\tau}(\bm{g}_{t}-\bm{m}_{t})\\ \bm{w}_{t+1}=\bm{w}_{t}-\eta_{t}\bm{m}_{t+1}.\end{cases}$
Equation (i) is SGD with momentum (14); notice how the SPP step size $\tau$ defines the momentum coefficient. Instead, (ii) corresponds to clipping, and here the SPP step size determines the clipping threshold. In other words, SGD with momentum is to online mean estimation what clipping is to online median estimation.
Connection to Error Feedback.
Methods such as (17) are also used in the distributed learning setting to improve communication and protect against corrupted nodes (Karimireddy et al., 2021). These methods are often called Error Feedback (EF) methods (Seide et al., 2014; Richtárik et al., 2021) and generally involve an update of the form $\bm{m}_{t+1}=\bm{m}_{t}+\mathcal{C}_{t}(\bm{g}_{t}-\bm{m}_{t})$ where $\mathcal{C}_{t}:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$ is a compression operator. Projections onto balls can be seen as a compression, since one can encode the resulting vector in fewer bits (Safaryan et al., 2021). In particular, centered clipping from (Karimireddy et al., 2021), and the Clip21 method (Khirirat et al., 2023) are using update (12) in the distributed setting, and are thus secretly estimating the geometric median of the gradients returned across all nodes.
SGD with standard clipping.
We will refer to (18) as clipped-SGD. The method has been studied extensively, often for the case $\beta=0$, see (Zhang et al., 2020a; Gorbunov et al., 2020; Koloskova et al., 2023; Puchkin et al., 2023) and references therein.
We present two motivations for using clipped-SGD. If $\beta=0$ then (18) is the same as (12) with cold start (as explained above). When $\beta\neq 0$, clipped-SGD can also be viewed as a SPP method for computing a robust mean estimator given by
The solution to (19) is $\mathbb{E}_{\bm{g}\sim\mathcal{G}}\left[\min\{1,\tfrac{c}{\|\bm{g}\|}\}\bm{g}\right]$, which is known as the Catoni-Giulini mean estimator (Catoni & Giulini, 2018; Lugosi & Mendelson, 2019). One step of SPP (9) applied to (19) gives (18) with $\beta=\tfrac{1}{1+\tau}$.
@@ 6 Experiments
We study both settings for synthetic data where we can control the level of heavy-tailedness, and for language modelling tasks with transformer models. In the following, we use VClip to refer to vectorwise clipping (12) and CClip for componentwise clipping  (13).
We consider the same three language modeling tasks as studied in (Kunstner et al., 2023): an encoder-only transformer for the PTB, a Transformer-XL model for the WikiText-2, and fine-tuning a DistillBERT for question-answering on the SQuAD dataset. We refer to Appendix C for details.
@@ 6.1 Gradient Estimation with Fixed Weights
In the first experiment, we verify the hypothesis that the choice of $\mathcal{D}$ matters for problem (4) – in particular, we show that estimating the median is more stable when the noise distribution is heavy-tailed.
We consider the problem of estimating a fixed vector $\bm{\hat{g}}$ from the oracle $\bm{g}\sim\bm{\hat{g}}+\bm{\xi}$ with $\bm{\xi}\sim\mathcal{P}$, under varying degree of heavy-tailedness of $\mathcal{P}$. For this purpose, we choose the standard, unskewed $\alpha$-stable distribution $\mathcal{P}_{\alpha}$.
Setup. We generate $\bm{\hat{g}}\in\mathbb{R}^{d}$ with $d=10$ where each component is generated i.i.d standard Gaussian. In each iteration, a sample $\bm{g}_{t}$ is generated as follows: each coordinate of $(\bm{g}_{t})_{i}$ is (independently) sampled from $\mathcal{P}_{\alpha}$ with location $(\hat{\bm{g}})_{i}$ and varying values for $0<\alpha\leq 2$. Importantly, the median of $\mathcal{P}_{\alpha}$ is equal to the location (i.e. $\bm{\hat{g}}$) for all values of $\alpha$. Recall that for $\alpha\leq 1$ the mean of $\mathcal{P}_{\alpha}$ is not defined, and otherwise equal to the median.
We run the SPP method (9) for $\mathcal{D}\in\{\tfrac{1}{2}\|\cdot\|^{2},\|\cdot\|_{2},\|\cdot\|_{1}\}$, with a fixed step size $\tau=0.01$ and track the relative $\ell_{2}$-error $\frac{\|\bm{m}_{t}-\bm{\hat{g}}\|_{2}}{\|\bm{\hat{g}}\|_{2}}$. The corresponding methods are called momentum, VClip, and CClip. For each method and distribution, we run $1000$ iterations and $50$ different seeds.
Discussion. From Fig. 3 we find that as the noise becomes more heavy-tailed (as $\alpha$ decreases), momentum fails to produce accurate estimates of $\bm{\hat{g}}$. On the other hand, both VClip and CClip are robust to heavy tails, as expected, as we showed that VClip and CClip are estimating the median.
@@
We also run momentum, VClip and CClip in order to estimate the full gradient for the language modeling tasks, where the weights are fixed at initialization. The setup is identical to the one described in (Kunstner et al., 2023), in particular Figure 1 therein. We present the results in Section C.1.
@@ 6.2 Training with Online Gradient Estimation
We consider the noise-perturbed least-squares problem $\min_{\bm{w}\in\mathbb{R}^{d}}\tfrac{1}{2}\|\bm{w}\|^{2}$ and use a gradient oracle given by $\bm{g}\sim\bm{w}+\bm{\xi}\in\mathbb{R}^{d}$, where $\bm{\xi}\sim\mathcal{P}$ is the noise term. We consider two setups:
@@
We run six different methods: on the one hand, the method depicted in Section 3, (5), where in each iteration we draw $n=5$ samples. We run three variations of (5), namely using the $\ell_{1}$- and $\ell_{2}$-sample median as well as the sample mean for $\bm{m}_{t}$. The $\ell_{2}$-median is approximately computed using the method proposed in (Vardi & Zhang, 2000).
@@
Figure 4: Least squares with heavy-tailed noise that is independent over components (left), state-dependent (middle), and where components of noise are dependent (right). All methods in grey use $n=5$ samples per iteration, all others only one. Note that for (S2), the sample mean immediately diverges, and hence does not appear in the plot. Shaded area depicts minimum and maximum value over 50 repetitions.
@@
Discussion. Fig. 4 shows that both $\ell_{1}$- and $\ell_{2}$-sample median attain the smallest objective, while the sample mean does not converge as predicted by our theory (see discussion after Proposition 3.3). From the online methods, CClip performs best in the scenario where the noise is componentwise independent, while VClip is better in the other case. We find that both VClip and CClip reach much smaller loss values than clipping the stochastic gradient directly, and thus clipping only the increment is better suited here. We also find that the $\ell_{1}$-sample median can deal with state-dependent noise (S2) as predicted by our theory.
Language modeling.
We now consider (17), that is training while simultaneously estimating the gradient, for the three language modelling tasks. We compare SGD-M, VClip and CClip, and clipped-SGD. We also add Adam as a baseline, but focus on the comparison among the SGD-type methods. For all methods, we tune the learning rate on a $\log_{10}$-scaled grid (reported in Table 1), displaying the one that attained smallest final training loss (averaged over $5$ seeds).
[…]
Discussion. Fig. 5 shows that VClip improves over SGD-M for PTB and WikiText-2, and is on par for SQuAD. CClip performs always worse, which might be due to the noise being not componentwise independent. However, VClip can not close the gap to Adam, and also performs worse than clipped-SGD for the WikiText-2 dataset.
@@ 7 Conclusion
We provide an analysis of SGD using the median of sampled gradients, and showed convergence guarantees under potentially heavy-tailed and state-dependent noise. In contrast, under the same assumption, SGD with the sample mean does not converge. We further provide a new approach to compute medians using the stochastic proximal point method. Our final algorithm proposes to use online estimates of the median gradient, requiring only one sample per iteration. We observe that this general framework intersects with many existing works on robust stochastic optimization, where the connection with median estimation was not known previously.
@@
The Appendix is organized as follows.
  • In Appendix A we provide missing proofs for the sample median method gradient descent (SMGD) presented in Section 3.
  • In Appendix B, we provide additional formulas and derivations for the problem of approximating the median, studied in Section 4.
  • In Appendix C, we provide additional plots and supplementary information on numerical experiments.
Appendix A Missing Proofs for Section 3
@@ A.1 On the Assumptions on the Sample Median
If Assumption 3.1 holds, and if $\ell$ is Lipschitz continuous, then without loss of generality we can assume that $\delta_{2}=\sigma_{2}=0$.
@@
In the following proof of Proposition 3.3 we will need an equivalent but more compact formulation of Assumption 3.1.
@@
We now verify that Assumption 3.1 holds true for the $\ell_{1}$-sample median, provided that the noise on the subgradient is linearly depending on an $\alpha$-stable distribution, and that the linear coefficients grow with the norm of the subgradient.
@@
Then, Assumption 3.1 holds with $\delta_{1}=\delta_{2}=0$, and $\sigma_{1}^{2}=C_{n}c_{1}$, $\sigma_{2}=C_{n}c_{2}$, for some constant $C_{n}=\mathcal{O}(1/n)$.
@@
where $=^{\mathrm{d}}$ denotes equality in distribution, $z$ is a standard symmetric $\alpha$-stable random variable in $\mathbb{R}$, and $\bm{\Sigma}(\bm{w})_{k}$ denotes the $k$-th row of $\bm{\Sigma}(\bm{w})$. Observe that from our Assumption 3.5 we can write that
@@
by setting $c_{1}^{\prime}=\rho_{\alpha}^{2}c_{1}$ and $c_{2}=\rho_{\alpha}^{2}c_{2}$, and where $\rho_{\alpha}>0$ is the best constant such that $\|\cdot\|_{\alpha}\leq\rho_{\alpha}\|\cdot\|_{2}$, which is $\rho_{\alpha}=d^{\tfrac{2-\alpha}{2\alpha}}$ is $\alpha\leq 2$, and $\rho_{\alpha}=1$ otherwise.
We start by computing the bias term that is required by Assumption 3.1.
@@
Now, let us compute $\mathbb{E}\left[\texttt{median}_{i\in[n]}\left(\xi^{(i)}_{k}\right)\right]$ for some $k\in\{1,\dots,d\}$. Denoting $r=(n-1)/2\in\mathbb{N}$ and by using the probability density function of the sample median (cf. (Maritz & Jarrett, 1978)), we have that:
@@
This shows that the first line of Assumption 3.1 holds with $\delta_{1}=\delta_{2}=0$.
@@
=\;= $\sum_{k=1}^{d}\mathbb{E}\left[\left(\texttt{median}_{i\in[n]}\left(\xi_{k}^{(i)}\right)\right)^{2}\right]$
=\;= $\sum_{k=1}^{d}\frac{n!}{r!r!}\mathbb{E}\left[\Bigl{(}\xi_{k}^{(i)}\Bigr{)}^{2}\Biggl{(}F_{\bm{w},k}\Bigl{(}\xi_{k}^{(i)}\Bigr{)}\Bigl{(}1-F_{\bm{w},k}\Bigl{(}\xi_{k}^{(i)}\Bigr{)}\Bigr{)}\Biggr{)}^{r}\right],$ (23)
where (23) follows from (Maritz & Jarrett, 1978, Equation 2.1). By using (21) with (22) and denoting the cdf of the standard symmetric $\alpha$-stable distribution by $F_{z}$, we further compute that:
@@
[…]
=:absent:=:\;= : $C_{n}\left(c_{1}^{\prime}+c_{2}^{\prime}\|\nabla\ell(\bm{w})\|^{2}\right).$
This shows that the second line of Assumption 3.1 holds with $\sigma_{1}^{2}=C_{n}\rho_{\alpha}^{2}c_{1}$ and $\sigma_{2}=C_{n}\rho_{\alpha}^{2}c_{2}$.
Finally we observe that, by again (Maritz & Jarrett, 1978, Equation 2.1), $C_{n}$ is the variance of the sample median of a set of i.i.d. standard symmetric $\alpha$-stable variables $\{z_{1},\dots,z_{n}\}$, where $z_{i}=^{\mathrm{d}}z$. As $n\to\infty$, it is well-known that the sample median is asymptotically normal with variance $1/(4nf^{2}(0))$, where $f$ denotes the pdf of $z$ (see (Maritz & Jarrett, 1978)). Hence we conclude that $C_{n}=\mathcal{O}(1/n)$ as $n\to\infty$. This concludes the proof. ∎
A.2 Rates for SMGD for Smooth Nonconvex Functions: Proof of Proposition 3.3
The following proposition is an application of biased SGD results to the SMGD algorithm with errors. It is possible to show that Assumptions 3.1 and 3.2 both imply Assumption 9 from (Demidovich et al., 2023). So from a qualitative point of view the result below is the same that what we would obtain by using Theorem 3 from (Demidovich et al., 2023). But doing so would produce a bound with worse constants, so we prefer to provide a proof exploiting directly our assumptions.
See 3.3
@@
In this proof we note $\mathbb{E}_{t}$ for the expectation taken conditionally to the filtration generated by $\bm{w}_{0},\dots\bm{w}_{t}$. In particular, we will have from our Assumption 3.2 on the errors together with the tower rule that $\mathbb{E}_{t}\|\bm{e}_{t}\|^{2}=\mathbb{E}_{t}\mathbb{E}\left[\|\bm{e}_{t}\|^{2}\ |\ \bm{g}_{t}^{(1)},\dots,\bm{g}_{t}^{(n)}\right]\leq\varepsilon^{2}$.
@@
where in the last inequality we used that (6) implies (20) (cf. Lemma A.2). Combine all the above inequalities to get

### v2

@@ 5 Theory for the Sample Median Gradient Method
Figure 1: First and second moment of sample median/mean for standard $\alpha$-stable distribution. We approximate $\mathbb{E}$ by running 10 00010\,000 trials. The sample median has much smaller variance (cf. (20)) compared to the sample mean.
@@ 6 Experiments
We study both settings for synthetic data where we can control the level of heavy-tailedness, as well as for language modelling tasks with transformer models. We consider the same three language modeling tasks as studied in (Kunstner et al., 2023): an encoder-only transformer for the PTB dataset, a Transformer-XL model for the WikiText-2 dataset, and fine-tuning a DistillBERT model for question-answering on the SQuAD dataset. We refer to Appendix˜C for details.
In the following, we use VClip to refer to vectorwise clipping (10) and CClip for componentwise clipping (11).
@@ 6.1 Gradient Estimation with Fixed Weights
In the first experiment, we verify the hypothesis that the choice of $\mathcal{D}$ matters for problem (8); in particular, we show that estimating the median is more stable when the noise distribution is heavy-tailed.
Setup. We consider the problem of estimating a fixed vector $\bm{\hat{g}}$ from the oracle $\bm{g}\sim\bm{\hat{g}}+\bm{\xi}$ with $\bm{\xi}\sim\mathcal{P}$, under varying degree of heavy-tailedness of $\mathcal{P}$. For this purpose, we choose the standard, unskewed $\alpha$-stable distribution $\mathcal{P}_{\alpha}$. We generate $\bm{\hat{g}}\in\mathbb{R}^{d}$ with $d=10$ where each component is generated i.i.d standard Gaussian. In each iteration, a sample $\bm{g}_{t}$ is generated as follows: each coordinate of $(\bm{g}_{t})_{i}$ is (independently) sampled from $\mathcal{P}_{\alpha}$ with location $(\hat{\bm{g}})_{i}$ and varying values for $0<\alpha\leq 2$. Importantly, the median of $\mathcal{P}_{\alpha}$ is equal to the location (i.e. $\bm{\hat{g}}$) for all values of $\alpha$. Recall that for $\alpha\leq 1$ the mean of $\mathcal{P}_{\alpha}$ is not defined, and otherwise equal to the median.
We run the SPP method (9) for $\mathcal{D}\in\{\tfrac{1}{2}\|\cdot\|^{2},\|\cdot\|_{2},\|\cdot\|_{1},H_{\mu}\}$, with a fixed step size $\tau=0.01$ and track the relative $\ell_{2}$-error $\frac{\|\bm{m}_{t}-\bm{\hat{g}}\|_{2}}{\|\bm{\hat{g}}\|_{2}}$. The corresponding methods are called momentum, VClip, CClip, and Huber, where we set $\mu=1.345$ according to Huber (1981). For each method and distribution, we run $1000$ iterations and $50$ different seeds.
Discussion. From Fig.˜2 we find that as the noise becomes more heavy-tailed (as $\alpha$ decreases), momentum fails to produce accurate estimates of $\bm{\hat{g}}$. On the other hand, both VClip and CClip are robust to heavy tails, as expected, as we showed that VClip and CClip are estimating the median. Notably CClip becomes more accurate (relatively) as $\alpha$ decreases. The Huber function produces results very similar to VClip, but slightly inferior. We refer to Fig.˜7 for additional ablations on the choice of $\tau$ and skewed noise.
@@
We also run momentum, VClip and CClip in order to estimate the full gradient for the language modeling tasks, where the weights are fixed at initialization. The setup is identical to the one described by Kunstner et al. (2023), in particular Figure 1 therein. We present the results in Section˜C.1.
@@ 6.2 Training with Online Gradient Estimation
We now present comparisons of various (online) gradient estimators when simultaneously training the weights, that is the distribution of gradients changes in each iteration.
@@
We consider the noise-perturbed least-squares problem $\min_{\bm{w}\in\mathbb{R}^{d}}\tfrac{1}{2}\|\bm{w}\|^{2}$ and use a gradient oracle given by $\bm{g}\sim\bm{w}+\bm{\xi}\in\mathbb{R}^{d}$, where $\bm{\xi}\sim\mathcal{P}$ is the noise term. We consider three setups:
@@
We run seven different methods: on the one hand, the method depicted in Section˜5, (18), where in each iteration we draw $n=5$ samples. We run three variations of (18), namely using the $\ell_{1}$- and $\ell_{2}$-sample median as well as the sample mean for $\bm{m}_{t}$. The $\ell_{2}$-median is approximately computed using the method proposed in Vardi & Zhang (2000).
@@
Figure 3: Least squares with heavy-tailed noise that is independent over components (left), state-dependent (middle), and where components of noise are dependent (right). All methods in gray use $n=5$ samples per iteration, all others only one. Note that for (S2), the sample mean immediately diverges, and hence does not appear in the plot. Shaded area depicts minimum and maximum value over 50 repetitions.
On the other hand, we run our online median estimator methods (15), with $\mathcal{D}\in\{\|\cdot\|_{2},~\|\cdot\|_{1},~H_{\mu}\}$, that is, VClip, CClip, and Huber. Importantly, these methods only receive one sample per iteration, and they do not involve any median computations (in contrast to $\ell_{1}$- and $\ell_{2}$-sample median). We also run clipped-SGD (cf. (16)) with $\beta=0.9$ and c=50c=50.444This choice is a tradeoff between slow convergence (large $c$) and high instability (small $c$). We run all methods with a learning rate $\eta_{t}=0.01$. For VClip, CClip, and Huber we set $\tau=1$, and we set again $\mu=1.345$ for Huber. We always run 50 repetitions and report averaged metrics.
Discussion. Fig.˜3 shows that both $\ell_{1}$- and $\ell_{2}$-sample median attain the smallest objective, while the sample mean does not converge as predicted by our theory (see discussion after Proposition˜5.3). From the online methods, CClip performs best in the scenario where the noise is componentwise independent, while VClip and Huber are better in the other case. Overall, Huber and VClip behave again very similarly (a more detailed analysis of Huber with different values of $\mu$ can be found in Fig.˜8). We find that both VClip and CClip reach much smaller loss values than clipping the stochastic gradient directly, and thus clipping only the increment is better suited here. We also find that the $\ell_{1}$-sample median can deal with state-dependent noise (S2) as predicted by our theory.
Language modeling.
We now consider the online methods (15), that is training while simultaneously estimating the gradient, for the three language modeling tasks. We compare SGD-M, VClip and CClip, and clipped-SGD. [^5] We also tried Huber (with $\mu=1.3$) on these problems, but did not observe an advantage over VClip (plots not shown). We also add Adam as a baseline, but focus on the comparison among the SGD-type methods. For all methods, we tune the learning rate on a $\log_{10}$-scaled grid (tuned values reported in Table˜1), displaying the one that attained smallest final training loss (averaged over $5$ seeds).
@@
Discussion. Fig.˜4 shows that VClip improves over SGD-M for PTB and WikiText-2, and is on par for SQuAD. CClip performs always worse, which might be due to the noise being not componentwise independent. However, VClip does not close the gap to Adam, and also performs worse than clipped-SGD for the WikiText-2 dataset.
@@ 7 Conclusion
We provide a new approach for robust gradient estimation using the stochastic proximal point method, with a focus on (online) estimators of the median gradient. We observe that this general framework intersects with many existing works on robust stochastic optimization, where the connection between clipping and median estimation was not known previously.
We provide an analysis of SGD using the median of sampled gradients (called SMGD), and show convergence guarantees under potentially heavy-tailed and state-dependent noise. In contrast, under the same assumption, SGD with the sample mean does not converge.
Finally, numerical experiments show that online versions of SMGD, even though not covered by this theory, perform equally well on problems with heavy-tailed noise, while methods based on mean estimation fail. However, when training transformer models on language tasks, this gap is far less pronounced; this could indicate that the level of heavy-tailedness in these applications is moderate.
Acknowledgments
We thank the Scientific Computing Core at the Flatiron Institute, a division of the Simons Foundation, for the compute facilities and support. U.Ş. and F.S. are partially supported by the French government under the management of Agence Nationale de la Recherche as part of the “Investissements d’avenir” program, reference ANR-19-P3IA-0001 (PRAIRIE 3IA Institute), and the European Research Council Starting Grant DYNASTY – 101039676. G.G. gratefully acknowledges the Flatiron Institute for their hospitality and financial support during the visit when this work was written.
@@ Acknowledgments
Appendix A Missing Proofs for Section˜5
@@ A.1 On the Assumptions on the Sample Median
If Assumption˜5.1 holds, and if $\ell$ is Lipschitz continuous, then without loss of generality we can assume that $\delta_{2}=\sigma_{2}=0$.
@@
In the following proof of Proposition˜5.3 we will need an equivalent but more compact formulation of Assumption˜5.1.
@@
We now verify that Assumption˜5.1 holds true for the $\ell_{1}$-sample median, provided that the noise on the subgradient is linearly depending on an $\alpha$-stable distribution, and that the linear coefficients grow with the norm of the subgradient.
@@
Then, Assumption˜5.1 holds with $\delta_{1}=\delta_{2}=0$, and $\sigma_{1}^{2}=C_{n}c_{1}$, $\sigma_{2}=C_{n}c_{2}$, for some constant $C_{n}=\mathcal{O}(1/n)$.
@@
where $^{\mathrm{d}}$ denotes equality in distribution, $z$ is a standard symmetric $\alpha$-stable random variable in $\mathbb{R}$, and $\bm{\Sigma}(\bm{w})_{k}$ denotes the $k$-th row of $\bm{\Sigma}(\bm{w})$. Observe that from Assumption˜5.5 we can write that
@@
by setting $c_{1}^{\prime}=\rho_{\alpha}^{2}c_{1}$ and $c_{2}=\rho_{\alpha}^{2}c_{2}$, and where $\rho_{\alpha}>0$ is the best constant such that $\|\cdot\|_{\alpha}\leq\rho_{\alpha}\|\cdot\|_{2}$, which is $\rho_{\alpha}=d^{\tfrac{2-\alpha}{2\alpha}}$ is for $\alpha\leq 2$, and $\rho_{\alpha}=1$ otherwise.
We start by computing the bias term that is required by Assumption˜5.1.
@@
Now, let us compute $\mathbb{E}\left[\texttt{median}_{i\in[n]}\left(\xi^{(i)}_{k}\right)\right$] for some $k\in\{1,\dots,d\}$. Denoting $r=(n-1)/2\in\mathbb{N}$ and by using the probability density function of the sample median (cf. Maritz & Jarrett (1978)), we have that:
@@
This shows that the first line of Assumption˜5.1 holds with $\delta_{1}=\delta_{2}=0$.
@@
$=\;$ $\sum_{k=1}^{d}\mathbb{E}\left[\left(\texttt{median}_{i\in[n]}\left(\xi_{k}^{(i)}\right)\right)^{2}\right]$
$=\;$ $\sum_{k=1}^{d}\frac{n!}{r!r!}\mathbb{E}\left[\Bigl(\xi_{k}^{(i)}\Bigr)^{2}\Biggl(F_{\bm{w},k}\Bigl(\xi_{k}^{(i)}\Bigr)\Bigl(1-F_{\bm{w},k}\Bigl(\xi_{k}^{(i)}\Bigr)\Bigr)\Biggr)^{r}\right],$ (24)
where (24) follows from Maritz & Jarrett (1978, Equation 2.1). By using (22) with (23) and denoting the cdf of the standard symmetric $\alpha$-stable distribution by $F_{z}$, we further compute that:
@@
$\leq\;$ $\frac{n!}{r!r!}\mathbb{E}\left[z^{2}\Biggl(F_{z}(z)(1-F_{z}(z)\Biggr)^{r}\right]\left(c_{1}^{\prime}+c_{2}^{\prime}\|\nabla\ell(\bm{w})\|^{2}\right)$
$=:\;$ $C_{n}\left(c_{1}^{\prime}+c_{2}^{\prime}\|\nabla\ell(\bm{w})\|^{2}\right).$
This shows that the second line of Assumption˜5.1 holds with $\sigma_{1}^{2}=C_{n}\rho_{\alpha}^{2}c_{1}$ and $\sigma_{2}=C_{n}\rho_{\alpha}^{2}c_{2}$.
Finally we observe that, again by Maritz & Jarrett (1978, Equation 2.1), $C_{n}$ is the variance of the sample median of a set of i.i.d. standard symmetric $\alpha$-stable variables $\{z_{1},\dots,z_{n}\}$, where $z_{i}=^{\mathrm{d}}z$. As $n\to\infty$, it is well-known that the sample median is asymptotically normal with variance $1/(4nf^{2}(0))$, where $f$ denotes the pdf of $z$ (see Maritz & Jarrett (1978)). Hence we conclude that $C_{n}=\mathcal{O}(1/n)$ as $n\to\infty$. This concludes the proof. ∎
A.2 Rates for SMGD for Smooth Nonconvex Functions: Proof of Proposition˜5.3
The following proposition is an application of biased SGD results to the SMGD algorithm with errors. It is possible to show that Assumptions˜5.1 and 5.2 both imply Assumption 9 from Demidovich et al. (2023). So from a qualitative point of view the result below is the same that what we would obtain by using Theorem 3 from Demidovich et al. (2023). But doing so would produce a bound with worse constants, so we prefer to provide a proof exploiting directly our assumptions.
See 5.3
@@
In this proof we note $\mathbb{E}_{t}$ for the expectation taken conditionally to the filtration generated by $\bm{w}_{0},\dots\bm{w}_{t}$. In particular, we will have from Assumption˜5.2 on the errors together with the tower rule that $\mathbb{E}_{t}\|\bm{e}_{t}\|^{2}=\mathbb{E}_{t}\mathbb{E}\left[\|\bm{e}_{t}\|^{2}\ |\ \bm{g}_{t}^{(1)},\dots,\bm{g}_{t}^{(n)}\right]\leq\varepsilon^{2}$.
@@
where in the last inequality we used that (19) implies (21) (cf. Lemma˜A.2). Combine all the above inequalities to get